OpenStudy (lncognlto):

How do I work out the absolute value for vector u if u = (1, -2, 6)?

OpenStudy (lncognlto):

I need to use it in the equation \[u.v = \left| u \right|\left| v \right| \cos \theta \]

OpenStudy (amistre64):

\[|u|=\sqrt{(u_x)^2+(u_y)^2+(u_z)^2}\] its just a distance formula

OpenStudy (amistre64):

square, add, sqrt

OpenStudy (lncognlto):

I remember now, yes, thanks again. :) This is work from the beginning of the year I am going over again.

OpenStudy (amistre64):

good luck :) you can also think of it as the sqrt of the dot product of u.u

OpenStudy (lncognlto):

Ok, cool, thanks :)